Calculate the pH of a solution containing 0.0334 M potassium hydrogen fumarate and 0.021 M dipotassium fumarate. The Ka values for fumaric acid are 9.50 × 10-4 (Ka1) and 3.30 × 10-5 (Ka2).
Answer #1

To do the calculations we have to consider the second ionization reaction as Potassium Hydrogen fumerate is formed due to the ionization of the 1st proton.

pKa2 = -logKa2 = -log (3.30 *10^-5) = 4.48

pH = pKa2 + log [dipotassium fumarate/Potassium hydrogen fumerate]

    = 4.48 + log [0.021/0.0334]

    =4.28